The scores of all pro golfers for a course are normally distributed with the mean of 70 and a standard deviation of 3. Suppose 36 pro golfers played the course today. Find the probability that the mean score of the 36 pro golfers exceeded 71.


1
Expert's answer
2022-03-08T21:35:36-0500

"P(\bar X>71)=P(Z>\frac{71-70}{\frac{3}{\sqrt{36}}})=P(Z>2)=1-P(Z<2)=0.0228."